OpenStudy (needshelpdfdf):

Tiles 24x – 6x2 = 0 2x(x + 3) = 0 Solution: x = 0, x = -3 14x – 7x2 = 0 6x(4 – x) = 0 Solution: x = 0, x = 4 2x2 + 6x = 0 x(4 – x) = 0 Solution: x = 0, x = 4 4x – x2 = 0 7x(2 – x) = 0 Solution: x = 0, x = 2

OpenStudy (alivejeremy):

24x – 6x2 = 0

OpenStudy (alivejeremy):

Wait.

OpenStudy (alivejeremy):

We match them right?

OpenStudy (needshelpdfdf):

yes

OpenStudy (alivejeremy):

Okay. :)

OpenStudy (alivejeremy):

24x – 6x2 = 0

OpenStudy (alivejeremy):

Is 6x(4 – x) = 0 Solution: x = 0, x = 4

OpenStudy (alivejeremy):

14x – 7x2 = 0

OpenStudy (alivejeremy):

is 7x(2 – x) = 0 Solution: x = 0, x = 2

OpenStudy (alivejeremy):

2x2 + 6x = 0

OpenStudy (alivejeremy):

is 2x(x + 3) = 0 Solution: x = 0, x = -3

OpenStudy (alivejeremy):

4x – x2 = 0

OpenStudy (alivejeremy):

is x(4 – x) = 0 Solution: x = 0, x = 4

OpenStudy (alivejeremy):

That's all ? right?

OpenStudy (mathmale):

Alivejeremy: Again I remind you that you are NOT to do the homework of others for them and that you are NOT to give out answers without involving the other person. Kindly respect OpenStudy guidelines.

OpenStudy (mathmale):

Note: x2 is incorrect if you mean "the square of x." Write x^2.

OpenStudy (mathmale):

Wrong. You are to guide the other person as that person tries to solve these problems himself. Matching or not, you are robbing the other person of the chance to learn skills and independence. Stop, please. Practice giving guidance...NOT answers.

OpenStudy (alivejeremy):

I'll tried.

OpenStudy (giitter):

Its ok. As he helped right? Maybe he forgot? He sometimes forgets stuff. Even though he did that it he never meant to. It sometimes is difficult for him. I understand that we couldn't give direct answers but I can help him with guidance. Also, don't delete his account because he's my best friend and I really do want to see him on here. If he allows me to help me with his improvement I will help him out! Do what you love, love what you do! ~Glitter

OpenStudy (alivejeremy):

The wolf has be summon

OpenStudy (alivejeremy):

been*

OpenStudy (giitter):

gtg